OpenStudy (anonymous):

4(2x − 1) − 3 = 9(x − 4)

OpenStudy (anonymous):

x=29

OpenStudy (anonymous):

8x-4-3=9x-36 8x-7=9x-36 subtract 8x from both sides -7=x-36 add 36 to both sides 29=x

OpenStudy (anonymous):

check it with 4[2(29)-1)-3=9(29-4).... both sides equal 225